OpenStudy (anonymous):

Help in Proving Identities: 2tanx/(1-tan^2x)+1/(2cos^2x-1) = cosx+sinx/(cosx-sinx) ? how ?

OpenStudy (primeralph):

Input question properly.

OpenStudy (anonymous):

that's the given.

OpenStudy (primeralph):

Can you draw the problem or use later.

OpenStudy (anonymous):

um i'll try new here in openstudy

OpenStudy (primeralph):

Okay, use the draw the button.

OpenStudy (anonymous):

\[\frac{ 2tanx }{ 1-\tan ^{2} x}+\frac{ 1 }{ 2\cos ^{2} x-1}=\frac{ cosx+sinx }{ cosx-sinx}\]

OpenStudy (anonymous):

sorry I took a long time doing it.

OpenStudy (anonymous):

how?

OpenStudy (primeralph):

Much better. Hold on.

OpenStudy (anonymous):

ok.

OpenStudy (primeralph):